A function f is given by f(4x-1) = 4x2+5. Evaluate f(15).
Answer: f(15) = Answer
905

Answer by greenestamps(13327)



905 would be the answer if the function were defined as f(x) = 4x^2+5:

f(15) = 4(15^2)+5 = 4(225)+5 = 900+5 = 905

But the function is defined as f(4x-1) = 4x^2+5. To evaluate f(15), you need to determine the value of x that makes 4x-1 equal to 15.

4x-1 = 15
4x = 16
x = 4

THAT is the value you need to substitute into the expression 4x^2+5 to get the answer.